# Long-Lived Electronic Coherences from First Principles

### v1

@@ III.1 Multiphoton excitation
To generate the initial superposition of states for subsequent TAB dynamical simulations, we explicitly model the multiphoton excitation of thiophene from its electronic ground state by the intense pulse employed in the experiment. To this end, we use TD-FOMO-CASCI, with the pulse modeled in the electric dipole approximation by adding an explicit time-dependent electric field to the Hamiltonian. Explicitly modeling the excitation process in this way implicitly includes many of the intricacies of strong-field multiphoton excitation: the bandwidth of the pulse (see fig. S1), the selection rules for multiphoton processes, any resonant enhancements, and the AC Stark effect. Here we have used a transform-limited pulse with parameters corresponding directly to the experiment: a frequency of 1.55 eV, field intensity of 1.5E+13 W/cm2, and pulse duration of 10 fs. Each of the 50 initial geometries (along with momenta to be used in the subsequent TAB simulations) were sampled from the zero-temperature vibrational Wigner distribution function, computed in the harmonic approximation based on frequencies calculated at the MP2 level. The direction of the linearly polarized pulse was randomized. The nuclei were frozen during our 40 fs long simulations, and the simulations were started 15 fs before the pulse is to reach maximum intensity. At the end of the simulations, any remaining ground state character was projected out of the CI wavefunction, and the remaining superposition was used as an initial condition for the subsequent TAB molecular dynamics simulations.
@@
State populations as a function of time are shown for a randomly selected representative example of such a simulation in Fig.2. Oscillations in state populations are observed to peak with the field intensity. After the field dies down, one is left with a superposition of excited states and any remaining ground state character. The final state populations depend not only on the molecular geometry but also on the polarization direction of the laser (See figs. S2-S4). In the present case, as in many of the trajectories, the population primarily settles in Rydberg states. The populations of high-lying valence states oscillate with the laser field as well (e.g., S'${}_{21}^{\textnormal{\textquotesingle}}$, $\pi_{2}\xrightarrow{}\pi_{1}^{*}$ in red), but these oscillations do not lead to significant population in the valence states at the end of the simulations.
@@
Figure 4: Time evolution of the superposition formed in Fig.2, visualized as positive (blue) and negative (red) electron density difference relative to the electronic ground state.
To gain better understanding of the resulting electronic coherence, we present a visualization in Fig.4. Specifically, we display the time evolution of the electronic density difference between the excited and the ground state for the same example simulation shown in Fig.2, which exists predominantly as a superposition of states S1, S3 and S9, with Rydberg 3s, 3pz, and 3dxz character, respectively. The snapshots cover 14.5 fs of TAB dynamics, which corresponds to two oscillation periods dictated by an energy gap of 0.6 eV between the S1 and S3 states. The coherent superposition results in a flow of electronic density (blue) from the upper side of the molecule to the lower one and back again. Additionally, interference with the 3dxz state introduces motion to appear in front and behind the molecular plane in between the two cycles. The negative part of the difference density (red) represents the hole from which the excited electron originates, i.e. the HOMO $\pi_{1}$ orbital. There is no noticeable evolution in the red part since both excitations share the same hole orbital. This is vital for experimental observation of coherences through ionization yield since a coherence would only be observable if the same final cationic state is accessible from all of the stationary states involved in the superposition.Merritt et al. (2021)
@@ III.2 TAB dynamics
$b(t,\Delta E)=\sum_{ijI}\lvert c_{i}^{I}(t)c_{j}^{I*}(t)\rvert\delta(E_{i}-E_{j}-\Delta E)P_{I},$ (13)
where $I$ indexes that 50 trajectories in the ensemble, $i$ and $j$ index electronic states, and $\delta$ is the Dirac delta function. Additionally, we weigh each trajectory contribution by the amount of excited state population after the laser pulse excitation, $P_{I}$, reflecting higher probability of occurrence. After a trajectory has collapsed to a single state, we no longer plot its signal to remove experimentally unobservable coherences, like those created following passage through conical intersections. Long-lasting coherences are observed as horizontal lines. Such coherences are observed at several gaps, the strongest of which can be seen at 0.6 eV. Collapses, triggered by the non-parallelity of the states (Eq. 7) or changes in state populations, result in the decay of these lines on the timescales ranging from tens of fs up to 100 fs. A video visualization of the coherent oscillations of the electronic wave packet and subsequent loss of coherence over time is included as supplementary material.
@@ III.3 Assignment of Coherences
Finally, we turn our attention to assigning the coherences observed in the experiment of Kaufman et al. In order to be consistent with the observed experimental data, several things must be true: a) both states involved in the coherence must be accessible upon multi-photon excitation; b) the PESs of the two states must be sufficiently parallel to support a long-lived coherence; c) the real-life gap between the states must correspond to the experimentally observed energy gap, 1.5 eV;Kaufman et al. (2023a) and d) for the coherence to be experimentally observable, the two states in question must correspond to the same cationic core.Merritt et al. (2021)
The above results strongly support the assignment of the experimentally observed coherences to two or more Rydberg state. First, we can eliminate the valence $\pi_{1}\xrightarrow{}\pi_{1}^{*}$ state because it is not readily excited by 4 photons (Fig. 3). The more accessible $\pi_{1}\xrightarrow{}\sigma_{1}^{*}$ state can be eliminated from consideration because excitation to this state results in ring opening, which destroys nuclear wave function overlap, thus killing the coherence very rapidly. Analysis of our simulated products reveals that 90 % of the molecules remain in the cyclical form at the end of the simulation while some of the molecules undergo ring-opening dissociation. This route has previously been studied in relation to the $\pi\xrightarrow{}\pi^{*}$ excitations which can undergo internal conversion to the $\pi\xrightarrow{}\sigma^{*}$ state, opening one of the C-S bonds.Prlj et al. (2015); Schnappinger et al. (2017) We also see involvement of these states in the ring-opening reaction.
@@ IV Conclusions
Herein, a novel combination of TDCI simulations of multiphoton absorption and TAB simulations of the subsequent molecular dynamics reveals long-lived coherences in thiophene. Results show the sole involvement of Rydberg states, with coherence half-lives of around 50 fs, allowing us to directly assign the long-lived coherences observed in recent ultrafast experiments.Kaufman et al. (2023b) Though we observe coherences between many pairs of states, based on several considerations, the experimentally observed signal can be assigned to coherence between 3s and 3d Rydberg states. The ability of the TAB approach to efficiently model dynamics on many electronic states and to discern between long- and short-lived coherences opens a range of opportunities in fields ranging from attoscience to quantum materials to photobiology.
@@ Acknowledgments
The authors acknowledge many valuable conversations with Tom Weinacht and Brian Kaufman. The authors gratefully acknowledge support from the National Science Foundation under grant NSF-1954519 and the Institute for Advanced Computational Science (IACS). J.S. acknowledges a postdoctoral fellowship from IACS.

### v2

@@ III.1 Multiphoton excitation
To generate the initial superposition of states for subsequent TAB dynamical simulations, we explicitly model the multiphoton excitation of thiophene from its electronic ground state by the intense pump pulse employed in the experiment. Our simulation of the pump is performed with frozen nuclei. We use TD-FOMO-CASCI to describe the many-electron dynamics, with the pulse modeled in the electric dipole approximation by adding an explicit time-dependent electric field to the Hamiltonian. Explicitly modeling the excitation process in this way implicitly includes many of the intricacies of strong-field multiphoton excitation: the bandwidth of the pulse (see fig. S1), the selection rules for multiphoton processes, any resonant enhancements, and the AC Stark effect. Here we have used a transform-limited pulse with parameters corresponding directly to the experiment: a frequency of 1.55 eV, field intensity of 1.5E+13 W/cm2, and pulse duration of 10 fs. Each of the 50 initial geometries (along with momenta to be used in the subsequent TAB simulations) were sampled from the zero-temperature vibrational Wigner distribution function, computed in the harmonic approximation based on frequencies calculated at the MP2 level. The direction of the linearly polarized pulse was randomized. The nuclei were frozen during our 40 fs long simulations to quantify the electronic response of each initial geometry to the field. The simulations were started 15 fs before the pulse is to reach maximum intensity. At the end of the simulations, any remaining ground state character was projected out of the CI wavefunctions, and the renormalized superpositions were used as initial conditions for the subsequent TAB molecular dynamics simulations.
@@
State populations as a function of time are shown for a single selected representative of such a simulation in Fig.2. Oscillations in state populations are observed to peak with the field intensity. After the field dies down, one is left with a superposition of excited states and any remaining ground state character. The final state populations depend not only on the molecular geometry but also on the polarization direction of the laser (See figs. S2-S4). In the present case, as in many of the trajectories, the population primarily settles in Rydberg states. The populations of high-lying valence states oscillate with the laser field as well (e.g., S'${}_{21}^{\textnormal{\textquotesingle}}$, $\pi_{2}\xrightarrow{}\pi_{1}^{*}$ in red), but these oscillations do not lead to significant population in the valence states at the end of the simulations.
@@
Figure 4: Time evolution of the superposition formed in Fig.2 during a single TAB trajectory. We visualize positive (blue) and negative (red) electron density difference relative to the electronic ground state.
To gain better understanding of the resulting electronic coherence, we present a visualization in Fig.4. Specifically, we display the TAB time evolution of the electronic density difference between the excited and the ground state for the selected initial superposition, which formation we have previously shown in Fig.2. It is predominantly a superposition of states S1, S3 and S9, with Rydberg 3s, 3pz, and 3dxz character, respectively. The snapshots cover 14.5 fs of TAB dynamics, which corresponds to two oscillation periods dictated by an energy gap of 0.6 eV between the S1 and S3 states. The coherent superposition results in a flow of electronic density (blue) from the upper side of the molecule to the lower one and back again. Additionally, interference with the 3dxz state introduces motion to appear in front and behind the molecular plane in between the two cycles. The negative part of the difference density (red) represents the hole from which the excited electron originates, i.e. the HOMO $\pi_{1}$ orbital. There is no noticeable evolution in the red part since both excitations share the same hole orbital. This is vital for experimental observation of coherences through ionization yield since a coherence would only be observable if the same final cationic state is accessible from all of the stationary states involved in the superposition.Merritt, Jacquemin, and Vacher (2021)
@@ III.2 TAB dynamics
$b(t,\Delta E)=\sum_{ijI}\lvert a_{i}^{I}(t)a_{j}^{I*}(t)\rvert\delta(E_{i}-E_{j}-\Delta E)P_{I},$ (14)
where $I$ indexes that 50 trajectories in the ensemble, $i$ and $j$ index electronic states, and $\delta$ is the Dirac delta function. Additionally, we weigh each trajectory contribution by the amount of excited state population after the laser pulse excitation, $P_{I}$, reflecting higher probability of occurrence. After a trajectory has collapsed to a single state, we no longer plot its signal to remove experimentally unobservable coherences, like those created following passage through conical intersections. Long-lasting coherences are observed as horizontal lines. Such coherences are observed at several gaps, the strongest of which can be seen at 0.6 eV. Collapses, triggered by the non-parallelity of the states (Eq. 7) or changes in state populations, result in the decay of these lines on the timescales ranging from tens of fs up to 100 fs. Note that collapses represent decoherence mechanism c), discussed in the intro—the loss of coherence is due to the decay of nuclear overlap between populations on different electronic states. A video visualization of the coherent oscillations of the electronic wave packet and subsequent loss of coherence over time is included as supplementary material.
@@ III.3 Assignment of Coherences
Finally, we turn our attention to assigning the coherences observed in the experiment of Kaufman et al. In order to be consistent with the observed experimental data, several things must be true: a) both states involved in the coherence must be accessible upon multi-photon excitation; b) the PESs of the two states must be sufficiently parallel to support a long-lived coherence; c) the real-life gap between the states must correspond to the experimentally observed energy gap, 1.5 eV;Kaufman et al. (2023a) and d) for the coherence to be experimentally observable, the two states in question must correspond to the same cationic core.Merritt, Jacquemin, and Vacher (2021)
The above results strongly support the assignment of the experimentally observed coherences to two or more Rydberg state. First, we can eliminate the valence $\pi_{1}\xrightarrow{}\pi_{1}^{*}$ state because it is not readily excited by 4 photons (Fig. 3). The more accessible $\pi_{1}\xrightarrow{}\sigma_{1}^{*}$ state can be eliminated from consideration because excitation to this state results in ring opening, which destroys nuclear wave function overlap, thus killing the coherence very rapidly. Analysis of our simulated products reveals that 90 % of the molecules remain in the cyclical form at the end of the simulation while some of the molecules undergo ring-opening dissociation. This route has previously been studied in relation to the $\pi\xrightarrow{}\pi^{*}$ excitations which can undergo internal conversion to the $\pi\xrightarrow{}\sigma^{*}$ state, opening one of the C-S bonds.Prlj, Curchod, and Corminboeuf (2015); Schnappinger et al. (2017) We also see involvement of these states in the ring-opening reaction.
@@
Symmetry has been noted to play a role in the observation of coherences created by ultrafast passage through conical intersections.Neville, Stolow, and Schuurman (2022) To within sampling error, all spectroscopic selection rules are naturally incorporated into our simulation of the pump pulse, as are the effects of symmetry on the subsequent dynamics. However, we have not explicitly computed the photoelectron yield. Thus, the degree to which these particular coherences modulate the photoelectron signal remains open for future work.
@@ IV Conclusions
Herein, a novel combination of TDCI simulations of multiphoton absorption and TAB simulations of the subsequent molecular dynamics reveals long-lived coherences in thiophene. Results show the sole involvement of Rydberg states, with coherence half-lives of around 50 fs, allowing us to directly assign the long-lived coherences observed in recent ultrafast experiments.Kaufman et al. (2023b) Though we observe coherences between many pairs of states, based on several considerations, the experimentally observed signal can be assigned to coherence between 3s and 3d Rydberg states. Here coherences between pairs of Rydberg states are long-lived because they share a common cationic core, and therefore nearly parallel PESs. The Rydberg electron itself orbits far enough from the nuclei that it exerts a relatively small force. Our simulations still underestimate the experimentally observed half-lives, which might be due to lack of recoherence. We would like to address this issue in the future by some type of collective post-processing of trajectories.S. Mattos, Mukherjee, and Barbatti (2024) It is worth noting that molecules in cavities may also achieve long-lived coherences, owing to dilution of the electron-phonon coupling in the resulting polariton state.Hu et al. (2022) The ability of the TAB approach to efficiently model dynamics on many electronic states and to discern between long- and short-lived coherences opens a range of opportunities in fields ranging from attoscience to quantum materials to photobiology.
@@ Acknowledgments
The authors acknowledge many valuable conversations with Tom Weinacht, Brian Kaufman, Ari Pereira, and Sam Vasquez. The authors gratefully acknowledge support from the National Science Foundation under grant NSF-1954519 and the Institute for Advanced Computational Science (IACS). J.S. acknowledges a postdoctoral fellowship from IACS.
Supplementary Material
See the supplementary material for following: details on multiphoton excitation calculations (pulse spectrum, initial superpositions, spherical sampling for different pulse energies), excited states properties, excited PESs along normal modes, basis set details, a video visualization of the coherent oscillations
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Jiří Suchan: Investigation (lead); Software (lead); Writing – original draft (equal); Writing – review & editing (equal). Benjamin G. Levine: Conceptualization (lead); Writing – original draft (equal); Writing – review & editing (equal).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.